### v1

@@ Abstract
[Abstract]Artificial intelligence (AI) is rapidly advancing in healthcare, enhancing the efficiency and effectiveness of services across various specialties, including cardiology, ophthalmology, dermatology, emergency medicine, etc. AI applications have significantly improved diagnostic accuracy, treatment personalization, and patient outcome predictions by leveraging technologies such as machine learning, neural networks, and natural language processing. However, these advancements also introduce substantial ethical and fairness challenges, particularly related to biases in data and algorithms. These biases can lead to disparities in healthcare delivery, affecting diagnostic accuracy and treatment outcomes across different demographic groups. This survey paper examines the integration of AI in healthcare, highlighting critical challenges related to bias and exploring strategies for mitigation. We emphasize the necessity of diverse datasets, fairness-aware algorithms, and regulatory frameworks to ensure equitable healthcare delivery. The paper concludes with recommendations for future research, advocating for interdisciplinary approaches, transparency in AI decision-making, and the development of innovative and inclusive AI applications.
[…]
keywords:
AI in Healthcare, Ethical Challenges, Bias Mitigation, Data Diversity, Fairness-aware Algorithms
articletype: Research articlejournal: Statistics in Medicinevolume: 00
1 Introduction
@@
Artificial intelligence (AI) is revolutionizing modern healthcare, dramatically transforming the ways we diagnose, treat, and manage diseases. The integration of AI into healthcare began in the late 20th century with systems like MYCIN 1 in the 1970s, which helped diagnose infections and recommend antibiotics, and CADUCEUS 2 in the 1980s, which emulated human diagnostic reasoning. These early systems laid the groundwork for today’s advanced machine learning and deep learning techniques, which now significantly enhance diagnostic accuracy, treatment personalization, and patient outcome predictions.
@@
As AI technologies have advanced, their impact on healthcare has grown exponentially. Modern AI applications, particularly deep learning, have enhanced image recognition, significantly improving diagnostic accuracy in fields such as radiology and pathology 3. Predictive analytics, powered by AI, are essential in patient monitoring and management, using real-time data to forecast potential patient deteriorations 4. Additionally, natural language processing (NLP) tools have revolutionized the handling of unstructured data, improving the functionality of electronic health record systems and facilitating more comprehensive patient care 5.
@@
Several key algorithms and technologies underpin these advancements. Neural networks, particularly convolutional neural networks (CNNs), are extensively used for medical image analysis, aiding in the detection and characterization of various pathological findings 6. Decision support systems incorporate diverse data, including genetic profiles and prior health records, to optimize treatment strategies 7. Among the notable AI tools, IBM Watson stands out for its application in cancer treatment, although its widespread adoption faces challenges 8.
@@
The integration of AI into healthcare brings substantial benefits, enhancing diagnostic and operational efficiencies while significantly reducing human error 9. AI’s ability to manage and analyze large datasets improves resource allocation and patient scheduling, directly impacting patient outcomes and satisfaction 10. In public health, AI’s predictive models are crucial for tracking disease patterns and managing epidemics, as evidenced during the COVID-19 pandemic 11. However, these advancements bring significant ethical and fairness challenges, particularly biases in training data that can lead to skewed AI models and disparities in healthcare outcomes among different demographic groups 12. For example, an algorithm used in US hospitals was biased against black patients in resource allocation 13, and dermatological AI showed lower diagnostic accuracy for conditions like melanoma in darker-skinned individuals due to training primarily on fair-skinned images  14. Similarly, AI tools for diagnosing depression have faced challenges when applied across different linguistic and cultural backgrounds because they were primarily trained on English-speaking, Western populations, leading to potential misdiagnoses in non-Western patients 15.
These examples highlight the urgent need for diverse datasets and transparent AI systems to ensure fairness and equity in healthcare delivery. Governments and regulatory bodies are responding to these challenges by crafting targeted policies and regulations. The European Union’s General Data Protection Regulation (GDPR) addresses privacy and ethical considerations in AI applications, setting a precedent for global data handling and protection standards 16. In the US, the Food and Drug Administration (FDA) is at the forefront of regulating medical AI applications, ensuring that new technologies are both safe and effective for clinical use 17.
This survey paper addresses the following questions:
  • How does bias manifest in AI systems used in healthcare?
  • What are the sources of bias?
  • What are the potential consequences of biased AI systems?
  • How do we address and mitigate these biases?
We investigate existing research on AI integration in healthcare, focusing on the emergence and origins of biases, their impact on outcomes, and strategies for detecting and mitigating these biases. The paper also delves into ethical and legal considerations related to AI in healthcare.
@@
This paper, Section 2, discusses the applications of AI in various healthcare dimensions. The biases exhibited by those AI applications, their root causes, and other details have been discussed in Section 3. Section 4 explores various approaches to address and mitigate bias. Ethical and legal considerations are discussed in Section 5. Section 6 identifies research gaps and future directions for enhancing AI in healthcare. This paper is concluded in Section 7.
@@ 2 Applications of AI in Healthcare
AI has emerged as a transformative force in healthcare, revolutionizing various medical domains through advanced algorithms, data analytics, and machine learning techniques. AI’s ability to process vast amounts of data with precision and speed has opened new avenues for enhancing diagnostic accuracy, personalizing treatments, and improving patient outcomes. This section explores the diverse applications of AI across several key areas of healthcare, highlighting its impact and potential for future innovations.
@@ 2.1 Cardiology
AI has significantly transformed cardiology by enhancing diagnostic accuracy, personalizing treatment, and improving patient outcomes 76, 77. AI algorithms are adept at analyzing cardiovascular data to facilitate early detection and diagnosis of heart diseases, including arrhythmias, heart failure, and coronary artery disease 78, 79. For instance, machine learning models can interpret echocardiograms with high precision, often surpassing human experts in diagnostic speed and accuracy 80. AI also assists in managing cardiovascular risk by integrating diverse patient data to predict individual risk factors more effectively 81. Furthermore, AI plays a crucial role in developing personalized treatment plans by analyzing patient data and predicting responses to various treatment modalities, optimizing therapeutic decisions 82. The integration of AI in clinical practice necessitates rigorous validation and ethical alignment to ensure patient safety and data security 83. The continuous evolution of AI technologies promises further innovations in diagnostic tools and therapeutic strategies, enhancing precision and predictive capabilities in whole person care whereby patients may be prescribed lifestyle changes specific to their context as seamlessly as medications tuned to their genomics 84. Moreover, AI can contribute to health equity by making advanced cardiac care accessible in underserved areas through telemedicine platforms and remote monitoring, thereby reducing disparities in cardiovascular health outcomes 85.
@@ 2.2 Ophthalmology
AI has revolutionized ophthalmology by enhancing diagnostic accuracy, improving patient outcomes, and streamlining clinical workflows. More and more, AI algorithms, especially deep learning models, are being used to look at complicated visual data from imaging methods like fundus photography and optical coherence tomography (OCT) 86. These models are very good at diagnosing common eye diseases like diabetic retinopathy, glaucoma, and age-related macular degeneration; they can sometimes match or beat the performance of experts 87. Automated analysis using AI not only speeds up the diagnostic process but also reduces human error, facilitating earlier and more precise interventions 88. Furthermore, AI’s predictive capabilities are being harnessed to forecast disease progression, which is crucial for conditions like glaucoma, where early detection can prevent severe vision loss 6. Using AI tools in clinical practice also includes using AI-driven decision support systems that help plan treatments by guessing how different types of treatments will work 89. Despite these advancements, challenges remain, including data privacy concerns, the need for large annotated datasets for training algorithms, and the integration of AI into existing clinical workflows 90. However, ongoing research and collaboration between AI technologists and ophthalmic experts are likely to overcome these obstacles, solidifying AI’s role in modern ophthalmology 91. This further opens the frontier for expansion of access to ophthalmological care in existing deserts through technology-enabled care models and non-specialist operators at the point of service, thereby addressing disparities in eye care availability and quality.
@@ 2.3 Dermatology
AI is transforming dermatology by enhancing diagnostic accuracy, personalizing treatment plans, and streamlining patient management. AI-powered tools, particularly CNNs, are highly effective in diagnosing skin cancer, matching or surpassing dermatologists in identifying melanomas 92, 93. Moreover, AI applications in dermatology extend beyond cancer detection. They include the assessment and management of chronic conditions such as psoriasis and atopic dermatitis, where AI algorithms help in monitoring disease progression and response to treatment 94. AI systems are also utilized in cosmetic dermatology, optimizing treatment recommendations based on individual facial analysis 95. Ensuring the efficacy and safety of AI tools requires rigorous validation processes with diverse datasets to mitigate bias and enhance generalizability 14. Integration into clinical workflows and user training are crucial to maximizing the benefits of AI in dermatological practice 96. As AI continues to evolve, continuous collaboration between technologists and clinicians is essential to address ethical considerations and improve patient outcomes in dermatology. Importantly, AI can democratize access to high-quality dermatological care by enabling remote consultations and diagnostics, thus reducing barriers for individuals in remote or underserved regions.
@@ 2.4 Neurology
AI has significantly advanced the field of neurology by enhancing diagnostic accuracy, personalizing treatment plans, and facilitating neurological research. AI techniques, particularly machine learning and deep learning, are pivotal in diagnosing neurological disorders such as Alzheimer’s disease and epilepsy by analyzing imaging data with high precision 97, 5. These technologies improve the interpretation of MRI and CT scans, often achieving higher accuracy than traditional methods 6. In treatment, AI algorithms personalize therapies based on patient data, optimizing treatment outcomes for diseases like Parkinson’s 98. Additionally, AI has revolutionized neuroprosthetics by creating adaptive interfaces, significantly improving the quality of life for patients with motor disabilities 99. Predictive analytics also play a crucial role in anticipating disease progression for conditions such as multiple sclerosis 100. AI’s application extends into research, facilitating the understanding of complex neurological phenomena and the development of innovative treatments 101. Overall, AI’s integration into neurology not only improves clinical practices but also opens new avenues to address rare diseases as well as more finely tune medications to the individual 102. Furthermore, AI can enhance health equity in neurology by enabling remote diagnosis and monitoring, ensuring that patients in underserved areas receive timely and accurate neurological 103.
@@ 2.5 Radiology and Cancer Treatments
AI is revolutionizing radiology and cancer treatment by enhancing imaging accuracy, enabling personalized therapy, and improving diagnostic workflows. AI algorithms, particularly deep learning models, have significantly improved the precision of image interpretation in radiology, aiding in the early detection and diagnosis of various cancers, such as breast and lung cancer 104. These models analyze imaging data much more rapidly than human radiologists, reducing diagnostic times and increasing efficiency 105. In cancer treatment, AI assists in formulating personalized treatment plans based on patient data and predictive analytics, which can predict treatment outcomes and suggest optimal therapies 106. AI applications in radiology extend to prognostic evaluations, predicting disease progression and survival rates, refining treatment protocols and developing follow-up strategies  107. These advancements are supported by large datasets and strict validation processes, ensuring reliability and clinical applicability 108.
@@ 2.6 Emergency Medicine and Critical Care
AI stands to significantly transform emergency medicine and critical care by enhancing patient triage, treatment efficacy, and time to diagnosis. As AI-powered tools like diagnostic algorithms have been incredibly good at the early identification of conditions like sepsis and its progression, there is the potential to improve patient outcomes 109, 110. The prediction of patient outcomes has also positioned AI to facilitate the optimization of critical resource allocation. During the COVID-19 pandemic, AI-based self-triage tools had a critical role to play in predicting cases and hospitalizations at the population level 111. AI-powered chatbots and virtual assistants are streamlining patient intake and symptom assessment 112, 113. Furthermore, AI’s role in triaging patients based on urgency has revolutionized care prioritization, improving ED operations and patient flow 114, 115. These advancements underscore AI’s potential to not only enhance emergency medical services but also pave the way for a more efficient, accurate, patient-centered healthcare system that extends care beyond the traditional brick-and-mortar facilities.

### v2

@@ Abstract
Artificial intelligence (AI) is rapidly advancing in healthcare, enhancing the efficiency and effectiveness of services across various specialties, including cardiology, ophthalmology, dermatology, emergency medicine, etc. AI applications have significantly improved diagnostic accuracy, treatment personalization, and patient outcome predictions by leveraging technologies such as machine learning, neural networks, and natural language processing. However, these advancements also introduce substantial ethical and fairness challenges, particularly related to biases in data and algorithms. These biases can lead to disparities in healthcare delivery, affecting diagnostic accuracy and treatment outcomes across different demographic groups. This review paper examines the integration of AI in healthcare, highlighting critical challenges related to bias and exploring strategies for mitigation. We emphasize the necessity of diverse datasets, fairness-aware algorithms, and regulatory frameworks to ensure equitable healthcare delivery. The paper concludes with recommendations for future research, advocating for interdisciplinary approaches, transparency in AI decision-making, and the development of innovative and inclusive AI applications.
Author summary
In this paper, we investigate the rapid advancement of artificial intelligence (AI) in healthcare, focusing on its role in improving efficiency and effectiveness across specialties such as cardiology, ophthalmology, and dermatology. We note that AI technologies enhance diagnostic accuracy, treatment personalization, and patient outcome predictions. However, these developments also pose significant ethical challenges, particularly concerning biases in data and algorithms that can create disparities in healthcare delivery. We examine the integration of AI in healthcare, highlighting the critical challenges related to bias and exploring strategies for mitigation. We emphasize the need for diverse datasets, fairness-aware algorithms, and regulatory frameworks to ensure equitable healthcare delivery. Our paper concludes with recommendations for future research, advocating for interdisciplinary approaches, transparency in AI decision-making, and the development of innovative and inclusive AI applications.
Introduction
@@
Artificial intelligence (AI) is revolutionizing modern healthcare, dramatically transforming the ways we diagnose, treat, and manage diseases. The integration of AI into healthcare began in the late 20th century with systems like MYCIN [1] in the 1970s, which helped diagnose infections and recommend antibiotics, and CADUCEUS [2] in the 1980s, which emulated human diagnostic reasoning. These early systems laid the groundwork for today’s advanced machine learning and deep learning techniques, which now significantly enhance diagnostic accuracy, treatment personalization, and patient outcome predictions.
@@
As AI technologies have advanced, their impact on healthcare has grown exponentially. Modern AI applications, particularly deep learning, have enhanced image recognition, significantly improving diagnostic accuracy in fields such as radiology and pathology [3]. Predictive analytics, powered by AI, are essential in patient monitoring and management, using real-time data to forecast potential patient deteriorations [4]. Additionally, natural language processing (NLP) tools have revolutionized the handling of unstructured data, improving the functionality of electronic health record systems and facilitating more comprehensive patient care [5].
@@
Several key algorithms and technologies underpin these advancements. Neural networks, particularly convolutional neural networks (CNNs), are extensively used for medical image analysis, aiding in the detection and characterization of various pathological findings [6]. Decision support systems incorporate diverse data, including genetic profiles and prior health records, to optimize treatment strategies [7]. Among the notable AI tools, IBM Watson stands out for its application in cancer treatment, although its widespread adoption faces challenges [8].
The integration of AI into healthcare has the potential to enhance diagnostic and operational efficiencies while assisting in reducing human error [9]. This is due to AI’s ability to manage and analyze large datasets, enabling more effective resource allocation and efficient patient scheduling, both of which can lead to improved patient outcomes and satisfaction [10]. In addition, AI-driven predictive models have proven useful in public health for tracking disease patterns and aiding in the management of epidemics, as seen during the COVID-19 pandemic [11]. These capabilities demonstrate AI’s potential to enhance healthcare delivery.
@@
However, alongside these advancements, substantial ethical and fairness challenges have emerged. Biases embedded within training data can lead to skewed AI models, resulting in disparities in healthcare outcomes across different demographic groups [12]. For instance, an algorithm used in US hospitals was biased against black patients in resource allocation [13], and dermatological AI showed lower diagnostic accuracy for conditions like melanoma in darker-skinned individuals due to training primarily on fair-skinned images  [14]. Similarly, AI tools for diagnosing depression have faced challenges when applied across different linguistic and cultural backgrounds because they were primarily trained on English-speaking, Western populations, leading to potential misdiagnoses in non-Western patients [15]. AI models trained predominantly on data from specific populations have exhibited lower diagnostic accuracy for underrepresented groups, exacerbating existing health disparities. These examples highlight the urgent need for diverse datasets and transparent AI systems to ensure fairness and equity in healthcare AI applications.
Efforts to address these challenges have increasingly gained attention from governments and regulatory bodies. Recognizing the potential of AI to both improve and exacerbate healthcare disparities, policymakers have taken initial steps toward creating safeguards. The European Union’s General Data Protection Regulation (GDPR), for example, provides a framework for ethical considerations in AI applications by addressing issues like data privacy and transparency [16]. In the United States, the Food and Drug Administration (FDA) has begun implementing guidelines to evaluate the safety and effectiveness of medical AI systems [17]. While these initiatives mark critical progress, they primarily focus on high-level governance and oversight, leaving significant gaps in addressing the technical challenges of bias detection and mitigation. Without comprehensive strategies that integrate policy, technical methodologies, and ethical considerations, these efforts risk being insufficient to tackle the systemic biases present in AI systems.
To bridge these gaps, this paper provides a significant contribution to the field of AI-driven healthcare by offering a comprehensive framework for understanding and addressing fairness and bias. It categorizes biases across the machine learning pipeline and aligns them with targeted detection and mitigation strategies. By integrating technical, ethical, and policy perspectives, the paper addresses key challenges in creating equitable AI systems and highlights actionable solutions. Through real-world examples from diverse healthcare domains, it bridges theoretical concepts with practical implementation. This work not only deepens the understanding of bias in healthcare AI but also provides researchers, practitioners, and policymakers with valuable tools and strategies for fostering fairness and inclusivity in AI applications.
@@
The paper is structured to help readers follow its key contributions. Section 1, discusses the applications of AI in various healthcare dimensions. The biases exhibited by those AI applications, their root causes, potential consequences, and fairness metrics and trade-offs have been discussed in Section 2. Section 3 explores various approaches to address and mitigate bias. Section 4 identifies research gaps and future directions for enhancing AI in healthcare. This paper is concluded in Section 5.
@@ 1 Applications of AI in Healthcare
AI has emerged as a transformative force in healthcare, leveraging advanced algorithms, data analytics, and machine learning techniques. By processing vast amounts of health data with remarkable speed and precision, AI systems are enhancing diagnostic accuracy, enabling personalized treatment plans, and ultimately elevating patient outcomes. This section explores the diverse applications of AI across several key areas of healthcare, highlighting its impact and potential for future innovations.
@@ 1.1 Cardiology
AI has the potential to transform cardiology by enhancing diagnostic accuracy, supporting personalized treatment, and contributing to improved patient care [18, 19]. Specifically, AI algorithms have proven effective in analyzing cardiovascular data to aid in the early detection and diagnosis of conditions such as arrhythmias, heart failure, and coronary artery disease [20, 21]. For instance, machine learning models have shown high precision in interpreting echocardiograms, at times surpassing human experts in diagnostic speed and accuracy [22]. Additionally, AI holds potential for managing cardiovascular risk by integrating diverse patient data to predict individual risk factors [23]. It also contributes to the development of personalized treatment plans by analyzing patient data and predicting responses to various treatment modalities, optimizing therapeutic decisions [24]. In oncology, AI tools like MesoNet have been used to predict patient survival by analyzing whole-slide digitized images. These tools have demonstrated greater accuracy in predicting patient survival compared to current pathology practices, offering new insights and potentially guiding better treatment decisions [25]. However, the integration of AI in clinical practice requires rigorous validation and careful ethical considerations to prioritize patient safety and uphold data security [26]. The continuous evolution of AI technologies holds promise for advancing diagnostic tools and therapeutic strategies, potentially improving precision and predictive capabilities in whole-person care. This could include recommending lifestyle changes tailored to individual contexts alongside medications adjusted to their genomic profiles [27]. Moreover, AI could play a role in supporting health equity by facilitating access to advanced cardiac care in underserved areas through telemedicine platforms and remote monitoring, which could help reduce disparities in cardiovascular health outcomes [28].
@@ 1.2 Ophthalmology
AI has demonstrated significant potential in ophthalmology by improving diagnostic accuracy for common eye diseases and streamlining clinical workflows. Deep learning models have been particularly effective in analyzing complex visual data from imaging techniques such as fundus photography and optical coherence tomography (OCT) [29]. These models are very good at diagnosing common eye diseases like diabetic retinopathy, glaucoma, and age-related macular degeneration, sometimes achieving expert-level sensitivity and specificity [30]. Automated analysis using AI not only speeds up the diagnostic process but also reduces human error, facilitating earlier and more precise interventions [31]. Furthermore, AI’s predictive capabilities are being harnessed to forecast disease progression, which is crucial for conditions like glaucoma, where early detection can prevent severe vision loss [6]. Using AI tools in clinical practice also includes using AI-driven decision support systems that help plan treatments by guessing how different types of treatments will work [32]. Despite these advancements, challenges remain, including data privacy concerns, the need for large annotated datasets for training algorithms, and the integration of AI into existing clinical workflows [33]. However, ongoing research and collaboration between AI technologists and ophthalmic experts are likely to overcome these obstacles, solidifying AI’s role in modern ophthalmology [34]. This further opens the frontier for the expansion of access to ophthalmological care in existing deserts through technology-enabled care models and non-specialist operators at the point of service, thereby addressing disparities in eye care availability and quality.
@@ 1.3 Dermatology
AI is advancing dermatology by improving diagnostic accuracy, personalizing treatments, and streamlining patient management. Deep learning models, such as CNNs, are effective in diagnosing skin cancer, demonstrating performance comparable to dermatologists in identifying melanomas and other skin conditions [35, 36]. Moreover, AI applications in dermatology extend beyond cancer detection to managing chronic conditions like psoriasis and atopic dermatitis. These algorithms assist in monitoring disease progression and treatment response, enhancing decision-making in telemedicine and clinical workflows [37]. AI systems are also utilized in cosmetic dermatology, optimizing treatment recommendations based on individual facial analysis [38]. Ensuring the efficacy and safety of AI tools requires rigorous validation processes with diverse datasets to mitigate bias and enhance generalizability [14]. Integration into clinical workflows and user training is crucial to maximizing the benefits of AI in dermatological practice [39]. As AI continues to evolve, continuous collaboration between technologists and clinicians is essential to address ethical considerations and improve patient outcomes in dermatology. Importantly, AI can democratize access to high-quality dermatological care by enabling remote consultations and diagnostics, thus reducing barriers for individuals in remote or underserved regions.
@@ 1.4 Neurology
AI has contributed to advancements in neurology by improving diagnostic accuracy and supporting research efforts. Techniques like machine learning and deep learning assist in analyzing imaging data for disorders such as Alzheimer’s and epilepsy [40, 5]. These technologies enhance the interpretation of MRI and CT scans, often achieving higher accuracy than traditional methods [6]. In treatment, AI algorithms personalize therapies based on patient data, optimizing outcomes for diseases like Parkinson’s [41]. Additionally, AI has advanced neuroprosthetics by enabling adaptive brain-machine interfaces (BMIs), which, when combined with physiotherapy, show potential to improve motor function in patients with severe motor disabilities [42]. Predictive analytics in healthcare highlight the importance of transparent algorithms for reliable validation and monitoring, which support applications in neurology, such as tracking disease progression in conditions like multiple sclerosis [43]. AI’s applications extend to research, facilitating the understanding of complex neurological phenomena and the development of innovative treatments [44]. For instance, StrokeSight, an EEG-based system, demonstrates how AI can improve neurology by providing fast, affordable stroke diagnoses and personalized treatments using deep learning [45]. Overall, AI’s integration into neurology not only enhances clinical practices but also opens new avenues to address rare diseases and develop more finely tuned medications tailored to individual patients [46].
@@ 1.5 Radiology and Cancer Treatments
AI is advancing radiology and cancer treatment by enhancing imaging accuracy, enabling personalized therapies, and improving diagnostic workflows. Deep learning algorithms have shown significant progress in interpreting medical images, supporting the early detection and diagnosis of cancers such as breast and lung cancer [47]. Machine learning models assist radiologists by identifying complex patterns in imaging data, aiding in faster and more efficient diagnoses [48]. In cancer treatment, AI assists in formulating personalized treatment plans based on patient data and predictive analytics, which can predict treatment outcomes and suggest optimal therapies [49]. AI applications in radiology extend to prognostic evaluations, predicting disease progression and survival rates, refining treatment protocols and developing follow-up strategies  [50]. The development of AI systems capable of outperforming human experts in tasks like breast cancer prediction highlights their potential in clinical settings. These advancements are supported by large datasets and strict validation processes, ensuring reliability and clinical applicability [51].
@@ 1.6 Emergency Medicine and Critical Care
AI stands to significantly transform emergency medicine and critical care by enhancing patient triage, treatment efficacy, and time to diagnosis. As AI-powered tools like diagnostic algorithms have been incredibly good at the early identification of conditions like sepsis and its progression, there is the potential to improve patient outcomes [52, 53]. The prediction of patient outcomes has also positioned AI to facilitate the optimization of critical resource allocation. During the COVID-19 pandemic, AI-based self-triage tools had a critical role to play in predicting cases and hospitalizations at the population level [54]. AI-powered chatbots and virtual assistants are streamlining patient intake and symptom assessment [55, 56]. Furthermore, AI’s role in triaging patients based on urgency has revolutionized care prioritization, improving ED operations and patient flow [57, 58]. These advancements underscore AI’s potential to not only enhance emergency medical services but also pave the way for a more efficient, accurate, patient-centered healthcare system that extends care beyond traditional brick-and-mortar facilities.